A. Suppose you deposit $1,100.00 into an account 4.00 years from today. Exactly 14.00 years from today the account is worth $1,756.00. What was the account's interest rate?

B. Suppose you need to have $53,880.00 in an account 17.00 years from today and that the account pays 8.00%. How much do you have to deposit into the account 9.00 years from today?

C. It is _______ for a company to issue equity than debt; it is ________ for an investor to buy equity in a company than debt in the same firm.

a. safer; safer

b. safer; riskier

c. riskier; safer

d. riskier; riskier

e. none of the above
Answer #1

Compound Amount Factor (CAF)=(F/P,i,N)=(1+i)^N

i=Annual Interest Rate , N=Number of Years

Future Value (F)=CAF*Present Value(P)

A.Present Value=P=$1100.00

Future Value =F=$1756.00

N=Number of years =(14-4)=10

1756=1100*CAF

CAF=1756/1100=1.596364

(1+i)^10=1.596364

1+i=1.596364^(1/10)=1.0479

Interest Rate =i=1.0479-1=0.0479

Interest Rate =4.79%

B.

B.Future Value =F=$53880

N=Number of years =(17-9)=8

i=Interest Rate=8%=0.08

CAE=(F/P,8%,8)=(1+0.08)^8

CAF=1.85093

F=1.85093*P

P=53880/1.85093=$29,109.69

C.

It is safer for a company to issue equity than debt; it is riskier for an investor to buy equity in a company than debt in the same firm.

ANSWER:

b. safer; riskier